# Deep learning-driven likelihood-free parameter inference for 21-cm forest observations

### v1

@@ Methods
Based on the above simulation, we propose a parameter inference method. The basic idea is to generate a large number of simulated data sets with relevant parameters and use these data sets to train a neural network called the NF 20 to approximate the posterior. By leveraging these simulated data sets, the trained network can quickly generate new posterior samples when observational data is obtained and tested. This bypasses the need to generate the 1D power spectrum at inference time, thereby amortizing expensive training costs across all future detections. The general method for building such models is called neural posterior estimation (NPE) 65.
In more detail, the NF is a method of probabilistic modeling using deep learning technology. Here, we use the notation of the INF for explanation. The differences between the INF and GNF, as shown in the Fig. 2, are merely the conditions and outputs. The core idea of NFs is to construct a reversible transformation $f_{P}$ based on the 1D power spectrum $P$, thereby obtaining the mapping relationship between a simple base distribution (e.g., a normal distribution) and a complex posterior distribution. It is a reversible neural network characterized by the ability to accurately calculate the Jacobian of the probability density function through network transformation. The NF can be mathematically expressed as 20
@@
where each $f_{P,~{}i}(u)$ represents a block of the NF. This iterative approach is central to NPE, where the goal is to train a parameter conditional distribution that approximates the true posterior distribution. This task translates into an optimization problem aimed at minimizing the expected Kullback-Leibler divergence (KLD) between the true and approximate posterior distributions 66:
@@
The training of the NF involves minimizing the loss function, which is the expected value of the cross-entropy between the true and model distributions 67:
@@
Our NF is neural spline flow (NSF) 68, specifically a rational-quadratic neural spline flow with autoregressive layers, RQ-NSF (AR). Each flow is connected by a series of autoregressive layers that share parameters. Specifically, the network uses the conditional and partial vector output from the previous flow (when $i$ is equal to 0, specifically the vector to be sampled from a standard distribution) as the input of a multilayer perceptron (MLP), and uses the output result as an RQ-NSF (AR). These parameters are defined as the width and height of $K$ bins, which ultimately form $K+1$ knots $\left\{\left(x^{(i)},y^{(i)}\right)\right\}_{i=0}^{K}$. The left and right endpoints are $(-B,-B)$ and $(B,B)$, respectively, and there are $K-1$ derivatives of knots $\left\{\dot{y}^{(i)}\right\}_{i=1}^{K-1}$, where the derivatives of the left and right endpoints are set to 1. Therefore, the data is divided into $K$ bins from $-B$ to $B$. Each stream is a map that maps $-B$ to $B$ and then $-B$ to $B$ again, so the y-axis is also divided into $K$ bins from $-B$ to $B$. The figure 1 in Ref. 68 defines the mapping relationship of the coupling transformation when $K$ is 10.
@@
NFs, unlike Bayesian methods, do not require an explicit definition of the likelihood function. It employs a likelihood-free approach, making it suitable for handling non-Gaussian data, where deriving an analytical likelihood is either computationally expensive or infeasible. Since the scale of the 21-cm forest is extremely small, generating the 1D power spectra of the 21-cm forest corresponding to a set of parameter points requires very high computational resources. This will make it difficult to obtain 1D power spectrum for any parameter, leading to incorrect estimates of the posterior distribution as presented in Supplementary Fig. 0.5. Therefore, our method consists of a GNF for generating the 1D power spectrum of arbitrary parameters and an INF for parameter inference. For the GNF, the conditions and outputs correspond to parameters and power spectra, respectively, whereas for the INF, the conditions and outputs correspond to power spectra and parameters, respectively. All the NF networks were trained using the “Xavier” 69 initialization for network parameters and the AdamW optimizer 70 on a single NVIDIA GeForce RTX A6000 GPU with 48 GB of memory.
The GNF is used to generate a 1D power spectrum corresponding to all parameters in the parameter space. Specifically, the GNF takes a parameter vector as input and finally outputs an estimate of the 1D power spectrum of the 21-cm forest. It selects $29\times 300$ from the black point data as the training set as presented in Supplementary Fig. 0.5, the blue point data as the verification set, and the red point as the test set. For GNFs, to prevent overfitting and achieve convergence, we employed early stopping with 15 epochs and utilized $L_{2}$ regularization 71. For both weaker and stronger heating effects, the activation function we used is rectified linear units (ReLU), the size of the hidden layers is 256, the number of bins is 8, the batch size is 256, the learning rate decay factor is 0.99, and $L_{2}$ regularization is set to 0.0001. The only difference is that for the weaker heating effect, the flow steps and transform blocks are both 8, and the learning rate is 0.0001, whereas for the stronger heating effect, the flow steps and transform blocks are both 6 and the learning rate is 0.0002.
The INF is designed to perform parameter reasoning based on data generated by the GNF. For the INF, its architecture is similar to the GNF. If the GNF can perfectly grasp the characteristics of the 1D power spectrum, then using the sampling points of the GNF within the prior range as the input of the INF, the INF can obtain appropriate inference results for the simulation data. For the INF, to ensure it adequately captures the posterior distribution of parameters, we employed 250 training epochs. For both weaker and stronger heating effects, the activation function we used is ReLU, the number of bins is 8, the flow steps and transform blocks are both 6, the batch size is 256, the learning rate decay factor is 0.99, and the learning rate is 0.0002. The only difference is that for the weaker heating effect, the size of the hidden layers is 128, whereas for the stronger heating effect, the size of the hidden layers is 256. Therefore, it receives the 1D power spectrum generated by the GNF as input during training to map it into a standard distribution. This enables simulation-based parameter estimation during testing. It selects the 1D power spectra generated by the GNF as the training set and trains for 300 epochs to ensure sufficient fitting. The remaining $300$ in the black point data is used as the test set as presented in Supplementary Fig. 0.5.
@@
We use the coefficient of determination of mean and standard deviation to test whether the 1D power spectrum generated by the GNF is reasonable, respectively. its expression is
@@
where $y_{\rm pred}$ and $y$ represent the mean value or standard deviation of the 1D power spectra generated by the GNF and the 1D power spectra corresponding to different scales generated by simulation, respectively. $\bar{y}$ represents the mean or standard deviation of the simulated generated 1D power spectra on different scales. To check whether the generated distribution under different conditions and different scales is consistent with the simulated situation, its JSD is also calculated by
@@
To evaluate the distribution characteristics of the 1D power spectra of the 21-cm forest on a specific scale under different astrophysical conditions, we used the skewness coefficient in statistical analysis as a key indicator to measure its Gaussianity. We calculated the skewness coefficient of the normalized 1D power spectrum at different scales $k$, which is defined as the cubic ratio of the third-order normalized moment to the standard deviation:
$S=E\left(\frac{P(k)-\mu}{\sigma}\right)^{3},$ (16)
where $P(k)$ represents the normalized 1D power spectrum at different scales $k$, $\mu$ is the mean value of $P(k)$, $\sigma$ is the standard deviation of $P(k)$, and $E$ represents the expected value. If the value of $S$ is close to zero, then it can be considered to be symmetrical on that scale and close to a Gaussian distribution. If $S$ is significantly different from zero, it indicates that the distribution has an obvious skew.
To quantify the correlation between different scales $k$ of the 1D power spectrum of the 21-cm forest, we use the Spearman correlation coefficient72. The Spearman correlation coefficient is a non-parametric measure of rank correlation that assesses the monotonic relationship between two variables. It does not assume a normal distribution of the data, making it more robust for handling non-Gaussian distributions. Given the significant non-Gaussian characteristics of our data, we chose the Spearman correlation coefficient to more accurately capture the correlations between different scales. This approach helps avoid errors that could arise from assuming a Gaussian distribution of the data. If all $n$ ranks are distinct integers, the Spearman correlation coefficient $\rho$ is defined as
@@
In our analysis, we employ the KS test 39 to assess the fit between the posterior distribution of the INF’s output and the theoretical distribution. The KS test compares the empirical distribution derived from the sample data with a specified theoretical distribution by measuring the maximum difference between their CDFs. This difference is known as the KS statistic $d$, which can be expressed as
@@
where $F_{n}(x)$ is the empirical CDF of the sample, $F(x)$ is the theoretical CDF. Theoretically, within a given confidence interval of the posterior distribution, the percentage of actual data should match the confidence level. Therefore, the theoretical CDF should be uniform. To interpret the d value in terms of statistical significance, we calculate a $p$-value using the distribution of the KS statistic under the null hypothesis. This $p$-value represents the probability of observing a KS statistic as extreme as $d$, assuming that the sample distribution is drawn from the theoretical distribution. The CDF of the KS statistic under the null hypothesis is given by
@@
All datasets are available from the corresponding authors upon reasonable request.
@@
This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).
Tian-Yang Sun performed most of the computation and wrote the majority of the manuscript. Yue Shao performed part of the computation and wrote part of the manuscript. Yichao Li and Yidong Xu wrote part of the manuscript. Xin Zhang and Yidong Xu proposed the study. Xin Zhang led the study and contributed to the manuscript writing. All authors discussed the results and commented on the manuscript.
@@
Figure 1: The SNR of the 1D power spectrum of the 21-cm forest corresponding to different heating levels in the WDM model with $m_{\rm WDM}$ = 6 keV. The blue, orange, green and red curves correspond to $f_{\rm X}$ = 0, $f_{\rm X}$ = 0.1, $f_{\rm X}$ = 1 and $f_{\rm X}$ = 3, respectively. The left panel presents the SNR with an integration time of 100 h on each source using SKA1-LOW, and the right panel presents the results for an integration time of 200 h on each source using SKA2-LOW. The error bars show the sample variances of the 1D power spectra of the 21-cm forest. The black dashed line in each panel indicates the SNR of 1.
@@
Figure 2: Network workflow chart (left panel) and schematic diagram of the architecture of the NF (right panel). Both the GNF and INF are composed of RQ-NSF (AR). In the GNF, the conditions are parameters, and the output is a 1D power spectrum. Conversely, in the INF, the condition is a 1D power spectrum, and the outputs are parameters. The network consists of a sequence of flows, each of which consists of a sequence of autoregressive layers that share parameters. Input the condition and the sampling vector obtained from the base distribution into the NFs to generate the output. The simulated power spectra data set is divided into a training set, a validation set, and a test set. The training set and validation set are used to train the GNF, which generates reasonable power spectra for any given parameters. The INF is trained using the 1D power spectra generated by the GNF to learn the parameter distribution corresponding to any given power spectrum. Different test sets of simulated power spectra are used to verify the rationality of the GNF and the INF, respectively.
@@
Figure 3: The mean value and $1\sigma$ distribution diagram of the GNF-generated 1D power spectra and simulated 1D power spectra corresponding to different $T_{\rm K}$ and $m_{\rm WDM}$ with an integration time of 100 h using SKA1-LOW. The blue area is the $1\sigma$ range of the 1D power spectrum generated by the GNF. The orange line and bar are the mean and $1\sigma$ error bar of the simulated power spectrum. $R_{\mu_{P(k)}}^{2}$ and $R_{\sigma_{P(k)}}^{2}$ are the correlation coefficients of the mean and standard deviation, respectively.
@@
Figure 4: $S$ of the 1D power spectrum at different integration time, $m_{\rm WDM}$, and $T_{\rm K}$. The blue, orange, and green curves in the left panel present the $S$ of the 1D power spectrum at different $k$ when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ = 60 K, and the SKA1-LOW integration time is 100 h, 1000 h, and infinity (pure signal), respectively. The blue, orange, and green curves in the middle panel present the $S$ of the 1D power spectrum at different $k$ when $T_{\rm K}$ = 60 K, and $m_{\rm WDM}$ is 3 keV, 6 keV, and 9 keV, respectively. The blue and orange curves in the right panel present the 1D power spectrum $S$ of the pure signal at different $k$ when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ is 60 K and 600 K, respectively.
Figure 5: The inference results for $m_{\rm WDM}$ and $T_{\rm K}$ of the INF. The blue line represents the most probable values under different parameters, and the blue area indicates the $1\sigma$ confidence interval. The value corresponding to each parameter is the average of the results of all test data under this parameter. The black diagonal line represents perfect (zero-error) results. The upper and lower panels present the constrained results of parameters using SKA1-LOW and SKA2-LOW, respectively. The left panels represent the predicted value $\xi_{\rm Pred}$ and the true value $\xi$ for $m_{\rm WDM}$, while the right panels show these values for $T_{\rm K}$.
@@
Figure 6: Constraints on $m_{\rm WDM}$ and $T_{\rm K}$ with the INF and Fisher matrix. The black contour is based on the INF and the blue contour is based on the Fisher matrix. The left panel presents the results when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ = 60 K with an integration time of 100 h using SKA1-LOW. The right panel presents the results when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ = 600 K with an integration time of 200 h using SKA2-LOW. Contours represent $1\sigma$ and $2\sigma$ confidence intervals.
@@
The resulting diagram of parameter inference using the NF directly without the INF for data enhancement. Due to the missing part of the 1D power spectrum corresponding to the parameters, the NF cannot obtain reasonable parameter inference results. The panel presents the results when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ = 60 K with an integration time of 100 h using SKA1-LOW. Contours represent $1\sigma$ and $2\sigma$ confidence intervals.
@@
JSDs between GNF and simulated power spectra at different $T_{\rm K}$ and $m_{\rm WDM}$. The left panel presents the average JSDs of the 1D power spectra generated by different parameters at different $k$ with an SKA1-LOW integration time of 100 h in the case of a weaker heating effect. The right panel presents the average JSDs of the power spectra generated by different parameters at different $k$ with an SKA2-LOW integration time of 200 h in the case of a stronger heating effect.
@@
The spearman correlation coefficient $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest at $z$ = 9 without considering thermal noise. The upper and lower panels present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with weaker ($T_{\rm K}$ = 60 K) and stronger ($T_{\rm K}$ = 600 K) heating effects, respectively. The left, middle, and right panels present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with $m_{\rm WDM}$ = 3 keV, $m_{\rm WDM}$ = 6 keV and $m_{\rm WDM}$ = 9 keV, respectively.
@@
The spearman correlation coefficient $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest at $z$ = 9 with thermal noise. The upper and lower panels present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with for weaker ($T_{\rm K}$ = 60 K with an integration time of 100 h using SKA1-LOW) and stronger ($T_{\rm K}$ = 600 K with an integration time of 200 h using SKA2-LOW) heating effects, respectively. The left, middle, and right panels present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with $m_{\rm WDM}$ = 3 keV, $m_{\rm WDM}$ = 6 keV and $m_{\rm WDM}$ = 9 keV, respectively.
@@
Precision distribution of parameter constrains for the INF for different $T_{\rm K}$ and $m_{\rm WDM}$. The upper left panel and lower left panel present the temperature constrained precision of the $T_{\rm K}$ using SKA1-LOW and SKA2-LOW, respectively. The upper and lower right panels present the constrained precision of $m_{\rm WDM}$ using SKA1-LOW and SKA2-LOW, respectively.

### v2

@@ Methods
Based on the above simulation, we propose a parameter inference method. The basic idea is to generate a large number of simulated data sets with relevant parameters and use these data sets to train a neural network, specifically an NF model 21 to approximate the posterior distribution. Compared to traditional generative models such as generative adversarial networks 72 and variational autoencoders 73, which often suffer from challenges like mode collapse, posterior collapse, and imprecise probability estimation 74, GNF provides a significant advantage by directly learning the exact likelihood function, which has been proven to enable high-quality data generation 22. In addition, NFs have demonstrated their gravitational wave parameter estimation capabilities by enabling real-time inference in high-dimensional spaces (up to 17 dimensions) 75, with scalability potential to higher dimensionality. Recent methodological advances reveal that emerging alternatives like diffusion models 76 achieve superior sample generation quality while addressing aforementioned limitations, yet their multi-step iterative processes incur substantial computational latency. By contrast, the NF framework maintains implementation practicality — a critical advantage requiring systematic evaluation when deploying such models in time-sensitive astrophysical applications. By leveraging these simulated data sets, the trained network can quickly generate new posterior samples when observational data is obtained and tested. This eliminates the need to generate the 1D power spectrum at inference time, effectively amortizing the computational cost of training across all future detections. The general method for building such models is called neural posterior estimation (NPE) 77.
In more detail, the NF is a method of probabilistic modeling using deep learning technology. Here, we use the notation of the INF for explanation. The differences between the INF and GNF, as shown in the Fig. 2, are merely the conditions and outputs. The core idea of NF is to construct a bijective transformation $f_{P}$ based on the 1D power spectrum $P$, thereby obtaining the mapping relationship between a simple base distribution (e.g., a normal distribution) and a complex posterior distribution. It is a type of reversible neural network characterized by the ability to accurately calculate the Jacobian determinant of the probability density function through network transformation. The NF can be mathematically expressed as 21
@@
where each $f_{P,~{}i}(u)$ represents a block of the NF. This iterative approach is central to NPE, where the goal is to train a parameter conditional distribution that approximates the true posterior distribution. This task translates into an optimization problem aimed at minimizing the expected Kullback-Leibler divergence (KLD) between the true and approximate posterior distributions 78:
@@
The training of the NF involves minimizing the loss function, which is the expected value of the cross-entropy between the true and model distributions 79:
@@
Our NF implementation is based on the neural spline flow (NSF) 80, specifically a rational-quadratic neural spline flow with autoregressive layers, RQ-NSF (AR). Each flow is connected by a series of autoregressive layers that share parameters. Specifically, the network uses the conditional and partially vectorized output from the previous flow (when $i$ is equal to 00, this is the vector sampled from a standard distribution) as the input of a multilayer perceptron (MLP), and the output of the MLP is used to parameterize the RQ-NSF (AR). These parameters are defined as the width and height of $K$ bins, which ultimately form $K+1$ knots $\left\{\left(x^{(i)},y^{(i)}\right)\right\}_{i=0}^{K}$. The left and right endpoints are $(-B,-B)$ and $(B,B)$, respectively, and there are $K-1$ derivatives of knots $\left\{\dot{y}^{(i)}\right\}_{i=1}^{K-1}$, where the derivatives of the left and right endpoints are set to 1. Therefore, the data is divided into $K$ bins from $-B$ to $B$. Each stream is a map that maps $-B$ to $B$ and then $-B$ to $B$ again, so the y-axis is also divided into $K$ bins from $-B$ to $B$. Figure 1 in Ref. 80 shows the mapping relationship of the coupling transformation when $K$ is $10$.
@@
NFs, unlike Bayesian methods, do not require an explicit definition of the likelihood function. They employ a likelihood-free approach, making it suitable for handling non-Gaussian data, where deriving an analytical likelihood is either computationally expensive or infeasible. Since the scale of the 21-cm forest data is extremely small, generating the 1D power spectra of the 21-cm forest corresponding to a set of parameter points requires even up to $40000$ core hours. This will make it difficult to obtain 1D power spectrum for any parameter, potentially leading to incorrect estimates of the posterior distribution as presented in Supplementary Fig. 0.8. Therefore, our method consists of a GNF for generating the 1D power spectrum of arbitrary parameters and an INF for parameter inference. The primary difference between these two is the conditioning: the GNF conditions on the parameters to generate power spectra, while the INF conditions on the power spectra to infer parameters. All the NF networks were trained using the “Xavier” 81 initialization for network parameters and the AdamW optimizer 82 on a single NVIDIA GeForce RTX A6000 GPU with 48 GB of memory.
The GNF is used to generate a 1D power spectrum corresponding to all parameters in the parameter space. Interpolation is a common method for filling in missing data and can serve as a supplementary approach for generating the 1D power spectra. However, as shown in Supplementary Fig. 0.8, due to the inherent cosmic variance and the correlations between different scales $k$, interpolation often results in lower $P(k)$ and less smooth transitions in the interpolated regions of parameter space. This highlights the limitations of interpolation methods, as they rely heavily on the assumption of smooth parameter variation, which is not always valid in the presence of cosmic variance. In contrast, the GNF model does not merely act as an interpolation tool but learns the underlying statistical relationships between parameters and the 1D power spectra, enabling it to reproduce both the smooth and stochastic features of the data. However, as is common in deep learning models, the GNF model may produce less reliable extrapolations in regions of parameter space that are underrepresented in the training data. For example, while the model performs well at high $T_{\rm K}$, it may yield unphysical results for large-scale modes at low $T_{\rm K}$. This limitation can be addressed by expanding the range of the training dataset to better cover the parameter space. Specifically, the GNF takes a parameter vector as input and finally outputs an estimate of the 1D power spectrum of the 21-cm forest. The training set is composed of $8700$ black points from the data, as shown in Supplementary Fig. 0.8, the blue point data as the validation set, and the red point as the testing set. For GNFs, to prevent overfitting and achieve convergence, we employed early stopping with a patience of 15 epochs and utilized $L_{1}$ regularization and $L_{2}$ regularization 83. A summary of key hyperparameters for GNF is provided in Supplementary Table 0.8.
The INF is designed to perform parameter inference based on data generated by the GNF. The architecture of the INF is similar to that of the GNF. Specifically, the INF receives the 1D power spectra generated by the GNF, conditioned on the 1D power spectra $P(k)$, and maps the corresponding parameters $\theta$ onto a standard distribution. By utilizing the bijective transformation property of the INF, this process enables the reverse prediction of the model parameters $\theta$. If the GNF can accurately capture the characteristics of the 1D power spectrum, then using the sampling points from the GNF within the prior range as the input of the INF, the INF can provide appropriate inference results for the simulation data. To prevent overfitting, the INF is trained on dynamically simulated data generated by the GNF in an iterative scheme. Specifically, during each training step of the INF, the GNF generates new samples based on the current training progress. These samples are used to update the INF’s parameters, and this iterative process continues until the loss function stabilizes, indicating that the training has converged. For this work, we use a batch size of $64$, and the GNF generates $128$ samples per batch during training. A summary of key hyperparameters for INF is provided in Supplementary Table 0.8. The remaining $300$ black points in the data are used as the testing set as presented in Supplementary Fig. 0.8.
@@
We use the coefficient of determination $R^{2}$ to evaluate whether the mean and standard deviation of the 1D power spectra generated by the GNF are reasonable. Its expression is:
@@
where $y_{\rm pred}$ and $y$ represent the mean or standard deviation of the 1D power spectra generated by the GNF and the corresponding values from the simulated 1D power spectra across different scales, respectively. $\bar{y}$ represents the overall mean or standard deviation of the simulated generated 1D power spectra on different scales. To check whether the distributions generated under different conditions and scales are consistent with the simulated situation, we also calculate JSD:
@@
To evaluate the distribution characteristics of the 1D power spectra of the 21-cm forest at specific scales under different astrophysical conditions, we used the skewness coefficient in statistical analysis as a key indicator to assess its Gaussianity. We calculated the skewness coefficient of the normalized 1D power spectrum at different scales $k$, which is defined as the ratio of the third standardized moment to the cube of the standard deviation:
$S=\mathbb{E}\left(\frac{P(k)-\mu}{\sigma}\right)^{3},$ (16)
where $P(k)$ represents the normalized 1D power spectrum at different scales $k$, $\mu$ is the mean value of $P(k)$, $\sigma$ is the standard deviation of $P(k)$, and $E$ represents the expected value. If the value of $S$ is close to zero, the distribution can be considered symmetric at that scale and close to a Gaussian distribution. If $S$ is significantly different from zero, it indicates that the distribution has a noticeable skewness.
To quantify the correlation between different scales $k$ of the 1D power spectrum of the 21-cm forest, we use the Spearman correlation coefficient84. The Spearman correlation coefficient is a non-parametric measure of rank correlation that assesses the monotonic relationship between two variables. It does not assume a normal distribution of the data, making it more robust for handling non-Gaussian distributions. Given the significant non-Gaussian characteristics of our data, we chose the Spearman correlation coefficient to more accurately capture the correlations between different scales. This approach helps avoid errors that could arise from assuming a Gaussian distribution of the data. If all $n$ ranks are distinct integers, the Spearman correlation coefficient $\rho$ is defined as
@@
In our analysis, we employ the KS test 43 to assess the fit between the posterior distribution of the INF’s output and the theoretical distribution. The KS test compares the empirical distribution derived from the sample data with a specified theoretical distribution by measuring the maximum difference between their CDFs. This difference is known as the KS statistic $d$, which can be expressed as
@@
where $F_{n}(x)$ is the empirical CDF of the sample, and $F(x)$ is the theoretical CDF. Theoretically, within a given confidence interval of the posterior distribution, the percentage of actual data should match the confidence level. Therefore, the theoretical CDF should be uniform. To interpret the d value in terms of statistical significance, we calculate a $p$-value using the distribution of the KS statistic under the null hypothesis. This $p$-value represents the probability of observing a KS statistic as extreme as $d$, assuming that the sample distribution is drawn from the theoretical distribution. The CDF of the KS statistic under the null hypothesis is given by
@@
The datasets for all figures in the text of this work are provided in Supplementary Data 1. Other datasets that support the findings of this work are available from the corresponding author upon reasonable request.
@@
This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).
Tian-Yang Sun performed most of the computation and wrote the majority of the manuscript. Yue Shao performed part of the computation and wrote part of the manuscript. Yichao Li, Yidong Xu and He Wang wrote part of the manuscript. Xin Zhang and Yidong Xu proposed the study. Xin Zhang led the study and contributed to the manuscript writing. All authors discussed the results and commented on the manuscript.
@@
Figure 1: The SNR of the 1D power spectrum of the 21-cm forest corresponding to different heating levels in the WDM model with $m_{\rm WDM}=6~{}{\rm keV}$. The blue, orange, green and red curves correspond to $f_{\rm X}=0$, $f_{\rm X}=0.1$, $f_{\rm X}=1$ and $f_{\rm X}=3$, respectively. a Depicts the SNR with an integration time of $100$ h on each source using SKA1-LOW. b Depicts the results for an integration time of $200$ h on each source using SKA2-LOW. The error bars show the sample variances of the 1D power spectra of the 21-cm forest. The black dashed line in each panel indicates the SNR of 1.
@@
Figure 2: Network workflow and schematic diagram of each part structure. a Depicts network workflow chart. b Depicts schematic diagram of the architecture of the NF. Both the GNF and INF are composed of RQ-NSF (AR). In the GNF, the conditions are parameters, and the output is a 1D power spectrum. Conversely, in the INF, the condition is a 1D power spectrum, and the outputs are parameters. The network consists of a sequence of flows, each of which consists of a sequence of autoregressive layers that share parameters. Input the condition and the sampling vector obtained from the base distribution into the NFs to generate the output. The simulated power spectra data set is divided into a training set, a validation set, and a testing set. The training set and validation set are used to train the GNF, which generates reasonable power spectra for any given parameters. The INF is trained using the 1D power spectra generated by the GNF to learn the parameter distribution corresponding to any given power spectrum. Different testing sets of simulated power spectra are used to verify the rationality of the GNF and the INF, respectively.
@@
Figure 3: The mean values and $1\sigma$ distribution of the GNF-generated 1D power spectra and simulated 1D power spectra. The data are derived from results of different $T_{\rm K}$ and $m_{\rm WDM}$, with an integration time of $100$ h using SKA1-LOW. a - c, d - f, g - i Present the 1D power spectra of 21-cm forest with $m_{\rm WDM}=3$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=9$ keV, respectively. The 1D power spectra of the 21-cm forest are presented in (a, d, g), (b, e, h) and (c, f, i) with $T_{\rm K}=47.5$ K, $T_{\rm K}=60$ K and $T_{\rm K}=82.5$ K, respectively. The blue shaded area is the $1\sigma$ range of the 1D power spectrum generated by the GNF. The orange line and bar are the mean and $1\sigma$ error bar of the simulated power spectrum. $R_{\mu_{P(k)}}^{2}$ and $R_{\sigma_{P(k)}}^{2}$ are the correlation coefficients of the mean and standard deviation, respectively.
@@
Figure 4: The skewness coefficient ($S$) of the 1D power spectrum at different integration time, $m_{\rm WDM}$, and $T_{\rm K}$. The blue, orange, and green curves in (a) present the $S$ of the 1D power spectrum at different $k$ when $m_{\rm WDM}=6$ keV and $T_{\rm K}=$ $60$ K, and the SKA1-LOW integration time is $100$ h, $1000$ h, and infinity (pure signal), respectively. The blue, orange, and green curves in (b) present the $S$ of the 1D power spectra at different $k$ when $T_{\rm K}=$ $60$ K, and $m_{\rm WDM}$ is $3$ keV, $6$ keV, and $9$ keV, respectively. The blue and orange curves in (c) present the 1D power spectrum $S$ of the pure signal at different $k$ when $m_{\rm WDM}=6$ keV and $T_{\rm K}$ is $60$ K and $600$ K, respectively.
Figure 5: P-P plot of the INF parameter inference. a Depicts the P-P plot of the INF inference with an integration time of $100$ h using SKA1-LOW and for a total of $2700$ power spectra with $m_{\rm WDM}$ ranging from $5$ keV to $7$ keV and $T_{\rm K}$ from $50$ K to $70$ K. b Depicts the P-P plot of the INF inference with an integration time of $200$ h using SKA2-LOW and for a total of $2700$ power spectra with $m_{\rm WDM}$ ranging from $5$ keV to $7$ keV and $T_{\rm K}$ from $400$ K to $800$ K. The gray shaded area is the range of the $3\sigma$ confidence interval. KS test $p$-values are denoted in the legend.
@@
Figure 6: The posterior distributions for $m_{\rm WDM}$ and $T_{\rm K}$ with the INF and Fisher matrix. The black contour is based on the INF and the blue contour is based on the Fisher matrix. a Depicts the results when $m_{\rm WDM}=6$ keV and $T_{\rm K}=$ $60$ K with an integration time of $100$ h using SKA1-LOW. b Depicts the results when $m_{\rm WDM}=6$ keV and $T_{\rm K}=600$ K with an integration time of $200$ h using SKA2-LOW. Contours represent $1\sigma$ and $2\sigma$ confidence intervals.
@@
The posterior distributions for $m_{\rm WDM}$ and $T_{\rm K}$ using the NF directly without the INF for data enhancement. Due to the missing part of the 1D power spectrum corresponding to the parameters, the NF cannot obtain reasonable parameter inference results. The panel presents the results when $m_{\rm WDM}=6$ keV and $T_{\rm K}=60$ K with an integration time of $100$ h using SKA1-LOW. Contours represent $1\sigma$ and $2\sigma$ confidence intervals.
@@
Expected 1D power spectrum of the 21-cm forest. A total of $100$ measurements of 10 comoving Mpc length segments along the line-of-sight neutral patch against 10 background light sources with $S_{150}=10$ mJy yielded the expected value for the 1D power spectrum of 21-cm forest at $z=9$. a, c Depict the 1D power spectra for different $m_{\rm WDM}$ values in the WDM models. The blue, orange, and green curves correspond to $m_{\rm WDM}=3$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=9$ keV, respectively. $T_{\rm K}$ in (a) and (c) are 60 K and 600 K, respectively. b, d Depict the 1D power spectra corresponding to the different $T_{\rm K}$. In all cases, $m_{\rm WDM}$ is set to $6$ keV. The blue, orange, and green curves in (b) correspond to $T_{\rm K}=40$ K, $T_{\rm K}=60$ K and $T_{\rm K}=80$ K. The blue, orange, and green curves in (d) to $T_{\rm K}=400$ K, $T_{\rm K}=600$ K and $T_{\rm K}=800$ K, respectively. The black solid and dashed lines are the thermal noises $P^{\mathrm{N}}$ expected for an integration time of $100$ h using SKA1-LOW and an integration time of $200$ h using SKA2-LOW, respectively. The error bars show sample variances of the 1D power spectra of the 21-cm forest.
@@
JSDs between GNF and simulated power spectra at different $T_{\rm K}$ and $m_{\rm WDM}$. a Depicts the average JSDs of the 1D power spectra generated by different parameters at different $k$ with an SKA1-LOW integration time of $100$ h in the case of a weaker heating effect. b Depicts the average JSDs of the power spectra generated by different parameters at different $k$ with an SKA2-LOW integration time of $200$ h in the case of a stronger heating effect.
@@
The spearman correlation coefficient $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest at $z=9$ without considering thermal noise. a - c, d - f Present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with weaker ($T_{\rm K}=60$ K) and stronger ($T_{\rm K}=600$ K) heating effects, respectively. $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest is presented in (a, d), (b, e) and (c, f) with $m_{\rm WDM}=3$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=9$ keV, respectively.
@@
The spearman correlation coefficient $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest at $z=9$ with thermal noise. a - c, d - f Present $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest with weaker ($T_{\rm K}=60$ K) and stronger ($T_{\rm K}=600$ K) heating effects, respectively. $\rho_{ij}$ estimated from the 1D power spectrum of 21-cm forest is presented in (a, d), (b, e) and (c, f) with $m_{\rm WDM}=3$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=9$ keV, respectively.